How much heat is needed to raise the temperature from 25oC to 50oC of 5kg

of substance? The specific heat capacity of the substance is 300J/kgoC


1
Expert's answer
2021-04-13T06:29:03-0400

The heat, needed to change the temperature of substance with specific heat capacity "c" by "\Delta T" degrees is "Q = c m \Delta T".

Hence, "Q = 300 \frac{J}{kg \cdot ^{\circ} C} \cdot 5 kg \cdot (50^{\circ} C - 25 ^{\circ} C) \approx 37.5 \cdot 10^3 J = 37.5 kJ".